Solve the equation: 3sqaure root -3x+4=2

1
Expert's answer
2021-03-09T12:35:51-0500

"Solution. \newline
\sqrt[3]{-3x+4}=2; \newline
-3x+4>0; \newline
-3x>-4; |*(-1) \newline
3x<4; \newline
x<4/3. \newline
\sqrt[3]{-3x+4}=2; |^3 \newline
-3x+4=8; \newline
-3x=4; \newline
x=-4/3. \newline
Answer: x=-4/3."